# Machine Learning GPU Graphics Processing Units

### v1

@@
Entropy-Based Uncertainty Modeling for Trajectory Prediction in Autonomous Driving
@@
Abstract
In autonomous driving, accurate motion prediction is essential for safe and efficient motion planning. To ensure safety, planners must rely on reliable uncertainty information about the predicted future behavior of surrounding agents, yet this aspect has received limited attention. This paper addresses the so-far neglected problem of uncertainty modeling in trajectory prediction. We adopt a holistic approach that focuses on uncertainty quantification, decomposition, and the influence of model composition. Our method is based on a theoretically grounded information-theoretic approach to measure uncertainty, allowing us to decompose total uncertainty into its aleatoric and epistemic components. We conduct extensive experiments on the nuScenes dataset to assess how different model architectures and configurations affect uncertainty quantification and model robustness.
11footnotetext: Bosch Center for Artificial Intelligence, Germany, Israel.22footnotetext: Department of Computer Science, University of Freiburg, Germany.
1 Introduction
In a machine learning driven Autonomous Driving (AD) stack, motion prediction connects the upstream task of environment perception with the downstream task of ego-motion planning (Hu et al., 2023). The role of a motion predictor is to infer the future motion of traffic agents relevant to the ego, ensuring safe and efficient progress toward a goal (Hagedorn et al., 2024). To achieve this, a predictor must tackle several challenges such as imperfect perception, complex interactions between traffic agents, as well as the multitude of distinct potential actions each agent might undertake, motivated by different goals. Such challenges drive the need to consider the problem in a probabilistic manner and incorporate uncertainty into prediction outputs. Accurately quantifying prediction uncertainty is essential for ensuring interpretability and building trust in the overall system.
In the AD community, the future motion of surrounding traffic agents is often modeled in the form of trajectories. Thus, probabilistic trajectory prediction involves capturing a distribution $p({y}|{x},\mathcal{D})$ of future trajectories ${y}$ conditioned on contextual data ${x}$ and a dataset $\mathcal{D}$. Contextual data ${x}$ usually contains past trajectories of surrounding agents and map information. There are different strategies for capturing this highly multi-modal distribution owing to the distinct actions or goals. Some methods attempt to directly predict the modes of the distribution along with their associated weights (Gao et al., 2020; Kim et al., 2021; Deo et al., 2022). Others use a parametric mixture distribution, such as a Gaussian Mixture Model (GMM), where the modes correspond to the predicted trajectories (Tolstaya et al., 2021; Varadarajan et al., 2022; Liu et al., 2024; Look et al., 2023). Alternatively, generative trajectory prediction models use well-known autoencoder or diffusion architectures to model latent variables and draw trajectory samples (Salzmann et al., 2020; Seff et al., 2023; Janjoš et al., 2023a; Jiang et al., 2023).
The majority of approaches for modeling the distribution of future trajectories, $p({y}|{x},\mathcal{D})$, in AD rely on neural networks. They are often underspecified by the available data, meaning that no single parameter configuration is favored. When considering uncertainty in the model parameters, the predictive distribution (Kapoor et al., 2022; MacKay, 1992) over future trajectories is computed as
@@ 1 Introduction
where ${\mathcal{W}}$ represents the neural network weights, and $p({\mathcal{W}}|\mathcal{D})$ represents the posterior distribution. The predictive distribution represents a Bayesian model average, meaning that instead of relying on a single hypothesis with a specific set of parameters, it considers all possible parameter configurations, weighted by their posterior $p({\mathcal{W}}|\mathcal{D})$. This marginalization process removes the reliance on a single weight configuration in the predictive distribution, resulting in better calibration and accuracy compared to traditional training methods (Wilson & Izmailov, 2020). Since the exact posterior is often intractable, various approximations $q({\mathcal{W}})$ have been developed, such as variational inference (Graves, 2011), Dropout (Gal & Ghahramani, 2016), Laplace approximation (Ritter et al., 2018), deep ensembles (Lakshminarayanan et al., 2017), or Markov Chain Monte Carlo (MCMC) methods (Welling & Teh, 2011).
Despite many successful approaches of approximating the posterior distribution, the AD prediction community has not yet tried to quantify or to decompose the uncertainty of trajectory prediction models in a theoretically principled manner (Wilson, 2020). This gap is notable, especially considering the potential benefits of decomposing the uncertainty. The total uncertainty can be decomposed into two types: aleatoric and epistemic uncertainty (Wimmer et al., 2023; Hüllermeier, 2021). Aleatoric uncertainty represents the inherent variability within the data, such as the equal likelihood of a vehicle turning left or right at a T-junction. This type of uncertainty cannot be reduced, even with more data. In contrast, epistemic uncertainty arises from the lack of knowledge or information and can be reduced by collecting more data (Wimmer et al., 2023). Knowledge of epistemic uncertainty is helpful in various contexts, e.g. risk-sensitive reinforcement learning (Depeweg et al., 2018) and Out-of-Distribution (OOD) detection (Amini et al., 2020). By understanding the sources of uncertainty, one can confirm that an autonomous vehicle finds itself in an OOD scenario by observing a higher epistemic uncertainty. This can be an important signal for a planner that uses predictions in its decision-making. By incorporating this information, planners can make more informed decisions and potentially take preventative actions in situations of high uncertainty.
@@
In this paper, we address the challenge of modeling the uncertainty of trajectory prediction models within the autonomous driving (AD) domain from a holistic perspective. We focus on the quantification and decomposition of uncertainty, as well as the influence of modeling choices related to the approximate posterior $q({\mathcal{W}})$. Our method employs an information-theoretic approach (Hüllermeier, 2021), which quantifies aleatoric uncertainty through conditional entropy and epistemic uncertainty using mutual information. Fig. 1 shows the predictive distribution $p(y|x,\mathcal{D})$ and its accompanying uncertainty values obtained by our proposed method for different scenarios of the nuScenes dataset (Caesar et al., 2020). We summarize our contributions as below:
@@
    We propose a novel method to quantify and decompose the uncertainty of trajectory prediction models, utilizing conditional entropy and mutual information to measure aleatoric and epistemic uncertainty. Both terms are approximated through Monte-Carlo (MC) sampling.
@@
    We analyze the relationship between uncertainty and prediction error in both in-distribution and out-of-distribution scenarios. Additionally, we evaluate robustness in handling out-of-distribution scenarios.
@@
    We study how modeling choices of the approximate posterior affect uncertainty calibration and model robustness, considering different configurations of deep ensembles and Dropout.
2 Measuring the Uncertainty for Trajectory Prediction
@@
This section details our method for decomposing uncertainty into aleatoric and epistemic components. We start by defining the problem of uncertainty decomposition in trajectory prediction in Sec. 2.1. Then, in Sec. 2.2, we describe our approach for calculating these uncertainties using an MC approximation. We provide pseudo-code in App. A. Finally, we discuss the limitations of our approach with possible avenues to address these limitations in Sec. 2.3.
2.1 Problem Statement
@@
Our method focuses on uncertainty quantification in trajectory prediction tasks. The problem is defined as predicting the future trajectory of a target agent in a driving scene based on current observations. Formally, let $x\in\mathbb{R}^{T_{in}\times F_{in}}$ represent the past features of an agent, where $T_{in}$ is the number of observed timesteps and $F_{in}$ denotes the number of input features, such as coordinates, velocities, accelerations, and other relevant data. In line with recent trajectory prediction literature (Deo et al., 2022; Liu et al., 2024; Kim et al., 2021), we also incorporate additional contexts, such as static map information and the past trajectories of surrounding agents, into the model input. A trajectory prediction model $f(x)=y$, parameterized by ${\mathcal{W}}$, uses this input to estimate a future trajectory $y\in\mathbb{R}^{T_{out}\times F_{out}}$. Here, $T_{out}$ represents the prediction horizon, and $F_{out}$ is the number of output features to predict, such as coordinates. Given the multi-modal nature of an agent’s future behavior, an extended version of this model predicts multiple future trajectories. The distribution over potential future outcomes, $p(y|x,{\mathcal{W}})$, can take various forms, such as a categorical distribution Deo et al. (2022), a mixture of Laplacians (Liu et al., 2024), a GMM (Nayakanti et al., 2023), or a non-parametric form (Jiang et al., 2023). Finally, we define an ensemble (Zhou, 2012) as a set of $M$ trajectory prediction models. These models may have different parameterizations and could belong to different model families. The ensemble can be constructed using various techniques, such as Dropout (Gal & Ghahramani, 2016), Stochastic Gradient Langevin Dynamics (SGLD) (Welling & Teh, 2011), or deep ensembles (Lakshminarayanan et al., 2017). This ensemble introduces a distribution $q({\mathcal{W}})$ over neural network parameters, which is an approximation to the true posterior $p({\mathcal{W}}|\mathcal{D})$ (Wilson & Izmailov, 2020).
@@
Our objective is to develop a method for uncertainty quantification to assess a model’s trustworthiness. However, the type of uncertainty to address is not always clear. On one hand, high uncertainty may stem from novel, previously unseen traffic scenarios. On the other hand, randomness arising from unpredictable driver behavior can lead to multiple plausible predictions. While previous works, such as Gilles et al. (2022) and Janjoš et al. (2023b), do not distinguish between uncertainty types, we argue that decomposing uncertainty is crucial for understanding the sources of a model’s predictions. Therefore, following concurrent literature (Der Kiureghian & Ditlevsen, 2009; Hüllermeier, 2021), we decompose uncertainty into epistemic and aleatoric components.
@@
2.2 Monte Carlo Approximation of the Conditional Entropy and Mutual Information as a Measure of Aleatoric and Epistemic Uncertainty
@@
To quantify uncertainty, we use entropy as a measure of total uncertainty. This allows us to frame our decomposition in terms of entropy subcomponents. Following Mobiny et al. (2021); Depeweg et al. (2018), we compute epistemic uncertainty as the difference between total and aleatoric uncertainty
@@
Above, ${\mathbf{I}}(y,{\mathcal{W}}|x,\mathcal{D})$ represents the mutual information between the model’s predictions and its parameters, while ${\mathbf{H}}(y|x,\mathcal{D})$ denotes the total entropy of the predictive distribution. The entropy of a distribution can be computed in closed form for simple cases, such as categorical distributions or univariate Gaussians. However, in trajectory prediction, the predictive distribution can take complex forms, such as a GMM (Nayakanti et al., 2023), making closed-form solutions to Eq. 2 unavailable. To address this, we use a Monte Carlo approximation. For a given input $x$, the entropy is approximated via set of $N$ samples from the predictive distribution, $y_{n}\sim p(y|x,\mathcal{D})$, as below
${{\mathbf{H}}}(y|x,\mathcal{D})=\mathbb{E}_{y}\left[-\log p(y|x,\mathcal{D})\right]\approx-\frac{1}{N}\sum_{n=1}^{N}{\log p(y_{n}|x,\mathcal{D})}=\hat{{\mathbf{H}}}(Y|x,\mathcal{D}).$ (3)
@@
$p(y|x,\mathcal{D})=\mathbb{E}_{p({\mathcal{W}}|\mathcal{D})}[p(y|x,{\mathcal{W}})]\approx\mathbb{E}_{q({\mathcal{W}})}[p(y|x,{\mathcal{W}})]=\frac{1}{M}\sum_{m=1}^{M}p(y|x,{\mathcal{W}}_{m}).$ (4)
The choice of the model composition $q({\mathcal{W}})$ significantly impacts the results, as different models may produce varied predictions, which will be explored further in Sec. 3. We then continue by inserting both Eq. 3 and 4 into the original problem as defined in Eq. 2
@@
Our proposed approach formalized in Eq. 25 assumes a generic form of the distribution $p(y|x,{\mathcal{W}}_{m})$. In practice, we use a continuous GMM that is ubiquitous in trajectory prediction for AD, see Sec. 4.1. Thus, we fit samples from a trajectory prediction model to a GMM, or directly use the GMM if the predictor provides one. Details around the GMM design choice can be found in App. B. In Fig. 2, we visualize GMMs fitted to the predictions from $M{=}3$ ensemble components, as well as samples from each GMM over a two-dimensional grid.

### v3

@@
Stochasticity in Motion: An Information-Theoretic
Approach to Trajectory Prediction
Aron Distelzweig1,2, Andreas Look1, Eitan Kosman1, Faris Janjoš1, Jörg Wagner1, Abhinav Valada2 1Bosch Center for Artificial Intelligence, Germany, Israel.2Department of Computer Science, University of Freiburg, Germany.
Abstract
In autonomous driving, accurate motion prediction is crucial for safe and efficient motion planning. To ensure safety, planners require reliable uncertainty estimates of the predicted behavior of surrounding agents, yet this aspect has received limited attention. In particular, decomposing uncertainty into its aleatoric and epistemic components is essential for distinguishing between inherent environmental randomness and model uncertainty, thereby enabling more robust and informed decision-making. This paper addresses the challenge of uncertainty modeling in trajectory prediction with a holistic approach that emphasizes uncertainty quantification, decomposition, and the impact of model composition. Our method, grounded in information theory, provides a theoretically principled way to measure uncertainty and decompose it into aleatoric and epistemic components. Unlike prior work, our approach is compatible with state-of-the-art motion predictors, allowing for broader applicability. We demonstrate its utility by conducting extensive experiments on the nuScenes dataset, which shows how different architectures and configurations influence uncertainty quantification and model robustness.
I Introduction
In a machine learning driven Autonomous Driving (AD) stack, motion prediction connects environment perception with ego motion planning [1]. The role of a motion predictor is to infer the future motion of relevant traffic agents to the ego agent, ensuring safe and efficient progress toward a goal [2]. To achieve this, a predictor must tackle several challenges, including imperfect perception, complex interactions between agents, and the multitude of potential actions that each agent could undertake. Addressing these challenges requires a probabilistic approach that incorporates uncertainty into prediction outputs, which is essential to ensure interpretability and build trust in the overall system.
In the AD community, the future motion of surrounding traffic agents is often modeled in the form of trajectories. Thus, probabilistic trajectory prediction involves capturing a distribution $p({y}|{x},\mathcal{D})$ of future trajectories ${y}$ conditioned on contextual data ${x}$ and a dataset $\mathcal{D}$. Contextual data ${x}$ usually contains past trajectories of surrounding agents and map information. There are different strategies for capturing this multi-modal distribution. Some methods attempt to directly predict the modes of the distribution along with their associated weights [3, 4, 5]. Others use a parametric mixture distribution, such as a Gaussian Mixture Model (GMM), where the modes correspond to the predicted trajectories [6, 7, 8, 9]. Alternatively, generative trajectory prediction models use well-known autoencoder or diffusion architectures to model latent variables and draw trajectory samples [10, 11, 12].
@@ I Introduction
Most approaches for modeling the distribution of future trajectories in AD rely on neural networks. They are often underspecified by the available data, meaning that no single parameter configuration is favored. When considering uncertainty in the model parameters, the predictive distribution $p({y}|{x},\mathcal{D})$ over future trajectories $y$ can be approximated [13] in the following manner
@@
where ${\mathcal{W}}$ represents the neural network weights and $p({\mathcal{W}}|\mathcal{D})$ represents the posterior distribution. The predictive distribution represents a Bayesian model average, meaning that instead of relying on a single hypothesis with a specific set of parameters, it considers all possible parameter configurations weighted by their posterior $p({\mathcal{W}}|\mathcal{D})$. This marginalization process removes the reliance on a single weight configuration in the predictive distribution, resulting in better calibration and accuracy [14]. Since the exact posterior is intractable, various approximations $q({\mathcal{W}})$ have been developed, such as variational inference [15], Dropout [16], Laplace approximation [17], deep ensembles [18], or Markov Chain Monte Carlo (MCMC) methods [19].
Despite numerous successful approaches to approximating the posterior distribution, the AD prediction community has yet to systematically quantify and decompose the uncertainty of trajectory prediction models in a theoretically principled manner [20]. A notable exception is [21], whose approach is inherently tied to predicting a categorical distribution over fixed trajectories, as done in [22] for example. This assumption limits their method, as state-of-the-art predictors do not generate their outputs by ranking a fixed set of trajectories. Overall, the literature gap is surprising given the importance of uncertainty modeling and its ubiquity in other domains [23, 24, 25].
In uncertainty modeling, total uncertainty can be categorized into two types: aleatoric and epistemic [26, 27]. Aleatoric uncertainty represents inherent variability in the data, such as the equal likelihood of a vehicle turning left or right at a T-junction. This type of uncertainty cannot be reduced, even with additional data. In contrast, epistemic uncertainty arises from a lack of knowledge and can be reduced by collecting more data [26]. Understanding epistemic uncertainty is valuable in various contexts, such as risk-sensitive reinforcement learning [23] and out-of-distribution (OOD) detection [28, 29]. By analyzing uncertainty sources, an Autonomous Vehicle (AV) can recognize OOD scenarios by detecting increased epistemic uncertainty. This can serve as a critical signal for a planner that relies on predictions for decision making. Incorporating this information enables planners to make more informed decisions and take precautionary actions in high-uncertainty situations. For example, a vehicle could signal for a human takeover in scenarios with high epistemic uncertainty [30]. Furthermore, uncertainty analysis can be applied not only to the behavior of other agents but also to the AV’s own planned trajectory, providing deeper insights into decision-making confidence and potential risks.
In this paper, we address the challenge of modeling the uncertainty of trajectory prediction models within the AD domain from a holistic perspective. We focus on the quantification and decomposition of uncertainty, as well as the influence of modeling choices related to the approximate posterior $q({\mathcal{W}})$. Our method employs an information-theoretic approach [27], which quantifies aleatoric uncertainty through conditional entropy and epistemic uncertainty using mutual information. Fig. 1 shows the predictive distribution $p(y|x,\mathcal{D})$ and its accompanying uncertainty values obtained by our proposed method for different scenarios of the nuScenes dataset [31]. We summarize our contributions as follows.
@@
    We propose a novel method to quantify and decompose the uncertainty of trajectory prediction models, utilizing conditional entropy and mutual information to measure aleatoric and epistemic uncertainty.
@@
    We analyze the relationship between uncertainty and prediction error in both in-distribution and out-of-distribution scenarios.
@@
    We study how posterior modeling choices impact uncertainty calibration and prediction robustness.
II Related Work
Anticipating the future motion of traffic participants is a critical component of autonomous driving systems [1]. Due to the safety-critical nature of these systems, it’s essential to account for uncertainties across the entire prediction stack. For instance, planners need to factor in motion prediction uncertainty to accurately assess the risks associated with various driving maneuvers [32]. In the following, we review related work on both motion prediction as well as uncertainty quantification and decomposition.
Motion Prediction for Autonomous Driving: The future motion of other traffic participants is influenced by a multitude of observable and unobservable factors, rendering it a challenging modeling task. These factors include, among others, the latent goals and preferences of traffic participants, social norms and traffic rules, complex interactions with surrounding traffic, as well as constraints induced by the static environment [33]. The shortcomings of a perception system that provides noisy and partial observations pose an additional challenge. These challenges necessitate a probabilistic formulation of the task to adequately model the uncertain and multi-modal nature of future motion. In general, prediction models typically consist of two components: a behavior backbone that encodes the traffic scene and a decoder that models the predictive distribution. We highlight various implementations of the two components below.
Early prediction approaches [22] propose encoding the past trajectory of observed traffic participants and the elements of the static environment (e.g., lane boundaries, crosswalks, traffic signs) by rendering the scene in a semantic bird’s eye view image and applying well-established convolutional neural networks. Such image-based representations of the scene have largely been replaced by vectorized representations [3, 4, 5, 34]. In a vectorized representation, all entities of the static and dynamic environment are approximated by a sequence of vectors. Models for sequential data, such as temporal convolutional networks [35] or recurrent neural networks are used to encode the sequences and interactions between entities are modeled using pooling operations, graph neural networks, or transformers.
The future motion of traffic participants is typically characterized by a sequence of states over multiple time steps, known as trajectories [36, 7]. Several strategies are employed to capture the highly multi-modal distribution over trajectories conditioned on the encoded scene. Many approaches represent the distribution by a set of trajectories with associated mode probabilities. The trajectories are either directly regressed by the model [37, 38, 4, 5] or fixed a beforehand [22]. In the case of predefined trajectories, the model is responsible for selecting the most likely ones. Other approaches use parametric mixture distributions, such as GMMs [39, 6, 7] or mixtures of Laplacians [8]. Alternatively, generative models such as conditional variational autoencoders [40, 41, 10, 11], generative adversarial networks [42, 43, 44], diffusion models [12], or normalizing flows [45] model the trajectory distribution via latent variables.
Uncertainty Modeling, Decomposition and Quantification: The majority of current trajectory prediction models solely account for aleatoric uncertainty by modeling a probability distribution over the output space [7]. To incorporate epistemic uncertainty in a theoretically sound manner, one can adopt a Bayesian framework [24, 23, 14, 20]. A Bayesian neural network assumes a distribution over the network weights instead of a point estimate to account for the lack of knowledge about the data generation process [27, 46]. Since analytically evaluating the posterior distribution over the weights is intractable for modern neural networks, approximate inference techniques such as Variational Inference (VI) or forms of MCMC must be considered [46]. Due to its simplicity, Monte-Carlo (MC) Dropout, which can be interpreted as an approximate VI method [16], is used by many perception approaches in AD [24, 25] and is also employed as one of two methods in [47] for modeling solely the epistemic uncertainty of a trajectory predictor. Another well-established approach to account for epistemic uncertainty are deep ensembles [18, 46, 14]. Prior work [32] uses deep ensembles to approximate the posterior distribution in their epistemic uncertainty-aware planning method. We apply MC Dropout as well as deep ensembles to approximate the uncertainty over network weights and systematically assess their performance in the context of trajectory prediction.
A common information-theoretical measure for the uncertainty is the entropy of the predictive distribution as a measure of the total uncertainty, which can be additively decomposed into the conditional entropy and mutual information, representing a measure of aleatoric and epistemic uncertainty [23, 27, 26]. Alternative measures based on variance are proposed in [23]. While variance-based measures are suitable in cases where the predictive distribution is a uni-modal Gaussian, it is less suitable for multi-modal outputs such as trajectories. Our approach thus relies on entropy-based measures to quantify the uncertainty of trajectory prediction models. However, variance can be useful in other contexts; [48] uses the variance of the predicted heat map over future positions as an uncertainty measure. Another variance-based uncertainty heuristic is proposed by [32] in the related field of motion planning for AD. This approach however only quantifies the epistemic uncertainty. Other methods learn proxy measures for the uncertainty of a trajectory prediction model without a proper decomposition: the approach in [49] trains separate models while [47] and [50] include additional heads with auxiliary tasks.
To the best of our knowledge, we are the first to offer a thorough and theoretically sound approach for modeling, decomposing, and quantifying uncertainties in trajectory prediction as a solid basis for future downstream applications. Existing approaches in the literature either fail to address all three aspects or rely on heuristics.
III Method
@@
This section details our method for decomposing the uncertainty into aleatoric and epistemic parts. We start by defining the problem of uncertainty decomposition in trajectory prediction in Sec. III-A. Then, in Sec. III-B, we describe our approach for approximating these uncertainties using a MC method. Finally, we discuss the limitations of our approach with possible avenues to address these in Sec. III-C.
III-A Problem Statement
@@
Our method focuses on uncertainty quantification in trajectory prediction tasks. The problem is defined as predicting the future trajectory of a target agent in a driving scene based on current observations. Formally, let $x\in\mathbb{R}^{T_{in}\times F_{in}}$ represent the past features of an agent, where $T_{in}$ is the number of observed timesteps and $F_{in}$ denotes the number of input features, such as coordinates, velocities, accelerations, and other relevant data. In line with recent trajectory prediction literature [5, 8, 4], we also incorporate additional context information, such as static map information and the past trajectories of surrounding agents, into the model input. A trajectory prediction model $f(x)=y$, parameterized by ${\mathcal{W}}$, uses this input to estimate a future trajectory $y\in\mathbb{R}^{T_{out}\times F_{out}}$. Here, $T_{out}$ represents the prediction horizon, and $F_{out}$ is the number of output features to predict, such as coordinates. Given the multi-modal nature of an agent’s future behavior, an extended version of this model predicts multiple future trajectories. The distribution over potential future outcomes, $p(y|x,{\mathcal{W}})$, can take various forms, such as a categorical distribution [5], a mixture of Laplacians [8], a GMM [34], or a non-parametric form [12]. Finally, we define an ensemble [51] as a set of $M$ trajectory prediction models. These models may have different parameterizations and could belong to different model families. The ensemble can be constructed using various techniques, such as Dropout [16], Stochastic Gradient Langevin Dynamics (SGLD) [19], or deep ensembles [18]. This ensemble introduces a distribution $q({\mathcal{W}})$ over neural network parameters, which is an approximation to the true posterior $p({\mathcal{W}}|\mathcal{D})$ [14].
@@
Our objective is to develop a method for uncertainty quantification to assess the trustworthiness of a model. However, the source of uncertainty is not always clear. On the one hand, high uncertainty may stem from novel, previously unseen traffic scenarios. On the other hand, randomness arising from unpredictable driver behavior can lead to multiple plausible predictions. While previous works such as [48] and [47] do not distinguish between uncertainty types, we argue that decomposing uncertainty is crucial for understanding the sources of potential error in prediction, which in turn supports safer and more effective downstream decision making. Therefore, following concurrent literature [52, 27], we decompose uncertainty into epistemic and aleatoric components.
@@
III-B Monte Carlo Approximation of the Conditional Entropy and Mutual Information as a Measure of Aleatoric and Epistemic Uncertainty
@@
In quantifying uncertainty, we use entropy as a measure of total uncertainty. This allows us to frame our decomposition in terms of entropy components. Following [53, 23], we compute epistemic uncertainty as the difference between total and aleatoric uncertainty
@@
Above, ${\mathbf{I}}(y,{\mathcal{W}}|x,\mathcal{D})$ represents the mutual information between the model’s predictions and its parameters, while ${\mathbf{H}}(y|x,\mathcal{D})$ denotes the total entropy of the predictive distribution. The entropy of a distribution can be computed in closed form for simple cases, such as categorical distributions or univariate Gaussians. However, in trajectory prediction, the predictive distribution can take complex forms, such as a GMM [34], making closed-form solutions to Eq. 2 unavailable. To address this, we use a Monte Carlo approximation. For a given input $x$, the entropy is approximated via set of $N$ samples from the predictive distribution, $y_{n}\sim p(y|x,\mathcal{D})$ as
${{\mathbf{H}}}(y|x,\mathcal{D})$ $=\mathbb{E}_{y}\left[-\log p(y|x,\mathcal{D})\right]$
$\approx-\frac{1}{N}\sum_{n=1}^{N}{\log p(y_{n}|x,\mathcal{D})}$
$=\hat{{\mathbf{H}}}(Y|x,\mathcal{D}).$ (3)
@@
$p(y|x,\mathcal{D})$ $=\mathbb{E}_{p({\mathcal{W}}|\mathcal{D})}[p(y|x,{\mathcal{W}})]$
$\approx\mathbb{E}_{q({\mathcal{W}})}[p(y|x,{\mathcal{W}})]$
$=\frac{1}{M}\sum_{m=1}^{M}p(y|x,{\mathcal{W}}_{m}).$ (4)
The choice of the model composition $q({\mathcal{W}})$ significantly impacts the results, as different models may produce varied predictions, which will be explored further in Sec. IV. We then continue by inserting both Eq. 3 and 4 into the original problem as defined in Eq. 2
@@
$\hskip 28.45274pt-\mathbb{E}_{q({\mathcal{W}})}\left[-\frac{1}{N}\sum_{n=1}^{N}{\log p(y_{n}|x,{\mathcal{W}})}\right]$
@@
Our proposed approach formalized in Eq. 25 assumes a generic form of the distribution $p(y|x,{\mathcal{W}}_{m})$. In practice, we use a continuous GMM that is ubiquitous in trajectory prediction for AD, see Sec. II. Thus, we fit samples from a trajectory prediction model to a GMM, or directly use the GMM if the predictor provides one. In Fig. 2, we visualize GMMs fitted to the predictions from $M{=}3$ ensemble components, as well as samples from each GMM over a two-dimensional grid.